OpenStudy (anonymous):

Simplify. The variable represents a nonnegative real number. √3x^5

OpenStudy (anonymous):

\[\sqrt{3x ^{5}}=\sqrt{3x ^{4}*x}=x ^{2}\sqrt{3x}\]